40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 12566     Entry time: Mon Oct 17 22:45:16 2016     In reply to: 12563     Reply to this: 12568
Author: gautam 
Type: Update 
Category: General 
Subject: AS beam centered on all OMs 

[ericq, lydia, gautam]

IMC realignment, Arm dither alignment

  • We started today by re-locking the PMC (required a c1psl restart), re-locking the IMC and then locking the arms
  • While trying to dither align the arms, I could only get the Y arm transmission to a maximum of ~0.09, while we are more used to something like 0.3 when the arm is well aligned this vent
  • As it turns out, Y arm was probably locked to an HOM, as a result of some minor drift in the ITMY optical table leveling due to the SOS tower aperture being left in over the weekend

ITMY chamber

  • We then resolved to start at the ITMY chamber, and re-confirm that the beam is indeed centered on the SRM by means of the above-mentioned aperture
  • Initially, there was considerable yaw misalignment on the aperture, probably due to the table level drifting because of the additional weight of the aperture
  • As soon as I removed the aperture, eric was able to re-dither-align the arms and their transmission went back up to the usual level of ~0.3 we are used to this vent
  • We quickly re-inserted the aperture and confirmed that the beam was indeed centered on the SRM
  • Then we removed the aperture from the chamber and set about inspecting the beam position on OM1
  • While the beam position wasn't terribly bad, we reasoned that we may as well do as good a job as we can now - so OM1 was moved ~0.5 in such that the beam through the SRM is now well centered on OM1 (see Attachment #1 for a CAD drawing of the ITMY table layout and the direction in which OM1 was moved)
  • Naturally this affected the beam position on OM2 - I re-centered the beam on OM2 by first coarsely rotating OM1 about the post it is mounted on, and then with the knobs on the mount. The beam is now well centered on OM2
  • We then went about checking the table leveling and found that the leveling had drifted substantially - I re-levelled the table by moving some of the weights around, but this has to be re-checked before closing up... 

BS/PRM chamber

  • The beam from OM2 was easily located in the BS/PRM chamber - it required minor yaw adjustment on OM2 to center the beam on OM3
  • Once the beam was centered on OM3, minor pitch and yaw adjustments on the OM3 mount were required to center the beam on OM4
  • The beam path from OM3 to OM4, and OM4 to the edge of the BS/PRM chamber towards the OMC chamber was checked. There is now good clearance (>2 beam diameters) between the beam from OM4 to the OMC chamber, and the green steering mirror in the path, which was one of the prime clipping candidates identified on Friday

OMC chamber

  • First, the beam was centered on OM5 by minor tweaking of the pitch and yaw knobs on OM4 (see Attachment #2)
  • Next, we set about removing the unused mirror just prior to the window on the AP table (see Attachment #3). PSL shutter was closed for this stage of work, in order to minimize the chance of staring directly into the input beam!
  • Unfortunately, we neglected checking the table leveling prior to removing the optic. A check after removing the optic suggested that the table wasn't level - this isn't so easy to check as the table is really crowded, and we can only really check near the edges of the table (see Attachment #3). But placing the level near the edge introduces an unknown amount of additional tilt due to its weight. We tried to minimize these effects by using the small spirit level, which confirmed that the table was indeed misaligned
  • To mitigate this, we placed a rectangular weight (clean) around the region where the removed mirror used to sit (see Attachment #3)Approximately half the block extends over the edge of the table, but it is bolted down. The leveling still isn't perfect - but we don't want to be too invasive on this table (see next bullet point). Since there are no suspended optics on this table, I think the leveling isn't as critical as on the other tables. We will take another pass at this tomorrow but I think we are in a good enough state right now. 
  • All this must have bumped the table quite a bit, because when we attempted re-locking the IMC, we noticed substantial misalignment. We should of course have anticipated this because the mirror launching the input beam into the IMC, and also MMT2 launching the beam into the arms, sits on this table! After exploring the alignment space of the IMC for a while, eric was able to re-lock the IMC and recover nominal transmission levels of ~1200 counts. 
  • We then re-locked the arms (needed some tip-tilt tweaking) and ran the dither again, setting us up for the final alignment onto OM6
  • OM5 pitch and yaw knobs were used to center the beam on OM6 - the resulting beam spot on OMPO-OMMTSM and OM6 are shown in Attachment #4 and Attachment #5 respectively. The centering on OMPO-OMMTSM isn't spectacular, but I wanted to avoid moving this optic if possible. Moreover, we don't really need the beam to follow this path (see last bullet in this section)
  • Beam path in the OMC chamber (OM5 --> OMPO-OMMTSM --> OM6 --> window was checked and no significant danger of clipping was found
  • Beam makes it cleanly through the window onto the AP table. We tweaked the pitch and yaw knobs on OM6 to center the beam on the first in-air pick off mirror steering the AS beam on the AP table. The beam is now visible on the camera, and looks clean, no hint of clipping
  • As a check, I wondered where the beam into the OMC is actually going. Turns out that as things stand, it is hitting the copper housing (see Attachment #6, it's had to get a good shot because of the crowded table...). While this isn't critical, perhaps we can avoid this extra scatter by dumping this beam?
  • Alternatively, we could just bypass OMPO-OMMTSM altogether - so rotate OM5 in-situ such that we steer the beam directly onto OM6. This way, we avoid throwing away half (?) the light in the AS beam. If this is the direction we want to take, it should be easy enough to make the change tomorrow

In summary...

  • AS beam has been centered on all steering optics (OM1 through OM6)
  • Table leveling has been checked on ITMY and OMC chambers - this will be re-checked prior to closing up
  • Green-scatter issue has to be investigated, should be fairly quick..
  • In the interest of neatness, we may want to install a couple of beam dumps - one to catch the back-reflection off the window in the OMC chamber, and the other for the beam going to the OMC (unless we decide to swivel OM5 and bypass the OMC section altogether, in which case the latter is superfluous)

C1SUSAUX re-booting

  • Not really related to this work, but we couldn't run the MC relief script due to c1susaux being unresponsive
  • I re-started c1susaux (taking care to follow the instructions in this elog to avoid getting ITMX stuck)
  • Afterwards, I was able to re-lock the IMC, recover nominal transmission of ~1200 counts. I then ran the MC relief servo
  • All shutters have been closed for the night
Attachment 1: OM1Moved.pdf  158 kB  | Hide | Hide all
OM1Moved.pdf
Attachment 2: IMG_3304.JPG  4.805 MB  | Hide | Hide all
IMG_3304.JPG
Attachment 3: OMCchamber.pdf  142 kB  | Hide | Hide all
OMCchamber.pdf
Attachment 4: IMG_3292.JPG  5.923 MB  | Hide | Hide all
IMG_3292.JPG
Attachment 5: IMG_3307.JPG  5.036 MB  | Hide | Hide all
IMG_3307.JPG
Attachment 6: IMG_3297.JPG  5.638 MB  | Hide | Hide all
IMG_3297.JPG
ELOG V3.1.3-